Click here to read the eBook: Market Value Ratios

PRICE/EARNINGS RATIO

A company has an EPS of $3.60, a book value per share of $37.08, and a market/book ratio of 2.5x. What is its P/E ratio? The stock price should be rounded to the nearest cent. Round your answer to two decimal places.

x
Answer #1

market/book ratio=market value/book value per share

market value=(2.5*37.08)=$92.7

PE ratio=market value/EPS

=92.7/3.6

=25.75